OpenStudy (anonymous):

the formula r = d/t gives the rate r for a distance d traveled in an amount of time t. A trip of 120 miles takes 2.5 hours. What was the rate for this trip? Please help! hate word problems!!

jimthompson5910 (jim_thompson5910):

The distance is d = 120 (given) and the time is t = 2.5 (also given) So plug those values into r = d/t to get your answer

OpenStudy (anonymous):

122.5?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

r = d/t r = 120/2.5 r = ???

jimthompson5910 (jim_thompson5910):

you added the two values by mistake

jimthompson5910 (jim_thompson5910):

when you should have divided them

OpenStudy (anonymous):

Ohhh!!

jimthompson5910 (jim_thompson5910):

yes, / is a fraction or division symbol (much like the \(\Large \div\) symbol)

OpenStudy (anonymous):

So its, 48?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Alright! Thanks!

jimthompson5910 (jim_thompson5910):

so 48 mph, yw